OpenStudy (anonymous):

How to calculate opportunity cost of moving to a new place? Like from A to B

OpenStudy (anonymous):

Mary is considering a choice between moving to New York City or continuing to live in Princeton, New Jersey. If she moved to New York City, the salary from her new job would be $190,000. Her rent for an acceptable apartment would be $6,000 per month and her apartment rental insurance premium would be $500 per month. At Princeton, she is making $150,000 from her job. Her apartment rent is $2,000 per month and her apartment rental insurance premium is $200 per month. What would be Mary’s opportunity cost of moving to New York? (Answer: $11,600) how comes $11,600! I don't get it

OpenStudy (anonymous):

To calculate Mary's opportunity cost of moving to New York, calculate Mary's net income in New York City for a year (after rent and insurance are paid) and compare it to her net income in Princeton, NJ for a year (after rent and insurance are paid). NYC: Salary - (rent*12) - (insurance*12) 190,000 - (6,000*12) - (500*12) = 190,000 - 72,000 - 6,000 = 112,000 Princeton, NJ: Salary - (rent*12) - (insurance*12) 150,000 - (2,000*12) - (200*12) = 150,000 - 24,000 - 2,400 = 123,600 Comparison: NYC - Princeton, NY 112,000 - 123,600 = -11,600 <--Opportunity cost So living in NYC would cost Mary 11,600 more than living in NJ.